# neurips_2024.bib

### v4

@@ 4.3 Hierarchical input and internal representation encoding
Training a conditional VAE requires [Sohn2015Learning] that the conditioning set $\bm{c}$ is concatenated to the input $\bm{x}$ before encoding occurs, and to the sample of latent variables $\bm{z}$ right before decoding. The same is also true, with the suitable adjustments, for any conditional generative approach where the target and the conditioning set must be processed jointly.
In order to ensure the usability and scalability of Carso across the widest range of input data and classifier models, we propose to perform such processing in a hierarchical and partially disjoint fashion between the input and the conditioning set. In principle, the encoding of $\bm{x}$ and $\bm{c}$ can be performed by two different and independent subnetworks, until some form of joint processing must occur. This allows to retain the overall architectural structure of the purifier, while having finer-grained control over the inductive biases [Mitchell1980Need] deemed the most suitable for the respective variables.
@@
In order to use the VAE decoder at inference time, the entire compression machinery for the conditioning set must be preserved after training, and used to encode the internal representations extracted. The equivalent input encoder may be discarded instead.
@@ 4.5 Robust aggregation strategy
Such choice produces a robust prediction much harder to take over in the event that an adversary selectively targets a specific input reconstruction. A heuristic justification for this property is given in Appendix D.
@@ 5 Experimental assessment
Experimental evaluation of our method is carried out in terms of robust and clean image classification accuracy within three different scenarios (a, b and c), determined by the specific classification task. The white-box threat model with a fixed $\ell_{\infty}$ norm bound is assumed throughout, as it generally constitutes the most demanding setup for adversarial defences.
@@ 5.1 Setup
The Cifar-10 [Krizhevsky2009LearningML] dataset is used in scenario (a), the Cifar-100 [Krizhevsky2009LearningML] dataset is used in scenario (b), whereas the TinyImageNet-200 [Chrabaszcz2017Downsampled] dataset is used in scenario (c).
@@
In scenarios (a) and (b), the classifier is trained according to [Cui2023Decoupled]; in scenario (c), according to [Wang2023Better]. Classifiers were always acquired as pre-trained models, using publicly available weights provided by the respective authors.
The purifier is trained on the VAE loss, using summed pixel-wise channel-wise binary cross-entropy as the reconstruction cost. Optimisation is performed by RAdam+Lookahead [Liu2020On, Zhang2019Lookahead] with a learning rate schedule that presents a linear warm-up, a plateau phase, and a linear annealing [Smith2017Cyclical]. To promote the learning of meaningful reconstructions during the initial phases of training, the KL divergence term in the VAE loss is suppressed for an initial number of epochs. Afterwards, it is linearly modulated up to its actual value, during a fixed number of epochs ($\beta$ increase) [Higgins2017Beta]. The initial and final epochs of such modulation are reported in Table 14.
@@
In each scenario, we report the clean and robust test-set accuracy – the latter by means of AutoAttack [Croce2020AutoAttack] – of the classifier and the corresponding Carso architecture.
For the classifier alone, the standard version of AutoAttack (AA) is used: i.e., the worst-case accuracy on a mixture of AutoPgd on the cross-entropy loss [Croce2020AutoAttack] with $100$ steps, AutoPgd on the difference of logits ratio loss [Croce2020AutoAttack] with $100$ steps, Fab [Croce2020Minimally] with $100$ steps, and the black-box Square attack [Andriushchenko2020Square] with $5000$ queries.
In the evaluation of the Carso architecture, the number of reconstructed samples per input is set to $8$, the logits are aggregated as explained in subsection 4.5, and the output class is finally selected as the $\operatorname*{\arg\max}$ of the aggregation. Due to the stochastic nature of the purifier, robust accuracy is assessed by a version of AutoAttack suitable for stochastic defences (randAA) – composed of AutoPgd on the cross-entropy and difference of logits ratio losses, across $20$ Expectation over Transformation (EoT) [Athalye2018Synthesizing] iterations with $100$ gradient ascent steps each.
@@
All experiments were performed on an NVIDIA DGX A100 system. Training in scenarios (a) and (c) was run on 8 NVIDIA A100 GPUs with 40 GB of dedicated memory each; in scenario (b) 4 of such devices were used. Elapsed real training time for the purifier in all scenarios is reported in Table 1.
Table 1: Elapsed real running time for training the purifier in the different scenarios considered.
@@ 5.2 Results and discussion
Table 2: Clean (results in italic) and adversarial (results in upright) accuracy for the different models and datasets used in the respective scenarios. The following abbreviations are used: Scen: scenario considered; AT/Cl: clean accuracy for the adversarially-pretrained model used as the classifier, when considered alone; C/Cl: clean accuracy for the Carso architecture; AT/AA: robust accuracy (by the means of AutoAttack) for the adversarially-pretrained model used as the classifier, when considered alone; C/randAA: robust accuracy for the Carso architecture, when attacked end-to-end by AutoAttack for randomised defences; Best AT/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial training alone (any model); Best P/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial purification (any model). Robust accuracies in round brackets are obtained using the Pgd+EoT [Lee2024Robust] pipeline, developed for diffusion-based purifiers. The best clean and robust accuracies per dataset are shown in bold. The clean accuracies for the models referred to in the Best columns are shown in Table 15 (in Appendix F).
Scen. Dataset AT/Cl C/Cl AT/AA C/rand-AA (Pgd+EoT) Best AT/AA Best P/AA (Pgd+EoT) (a) Cifar-10 0.9216 0.8686 0.6773 0.7613 (0.7689) 0.7107 0.7812 (0.6641) (b) Cifar-100 0.7385 0.6806 0.3918 0.6665 0.4267 0.4609 (c) TinyImageNet-200 0.6519 0.5632 0.3130 0.5356 0.3130
@@
Comparing the robust accuracy of the classifier model used in scenario (a) [Cui2023Decoupled] with that resulting from the inclusion of the same model in the Carso architecture, we observe a $+8.4\%$ increase. This is counterbalanced by a $-5.6\%$ clean accuracy toll. The same version of Carso further provides a $+5.03$ robustness increase w.r.t. the current best AT-trained model [Peng2023Robust] that employs a 3$\sim$ 3$\times$ 3 × larger RaWideResNet-70-16 model.
In addition, our method provides a remarkable $+9.72\%$ increase in robust accuracy w.r.t. to the best adversarial purification approach [Lin2024Robust], a diffusion-based purifier. However, the comparison is not as straightforward. In fact, the paper [Lin2024Robust] reports a robust accuracy of $78.12\%$ using AutoAttack on the gradients obtained via the adjoint method [Nie2022DiffPure]. As noted in [Lee2024Robust], such evaluation (which uses the version of AutoAttack that is unsuitable for stochastic defences) leads to a large overestimation of the robustness of diffusive purifiers. As suggested in [Lee2024Robust], the authors of [Lin2024Robust] re-evaluate the robust accuracy according to a more suitable pipeline (Pgd+EoT, whose hyperparameters are shown in Table 12), obtaining a much lower robust accuracy of $66.41\%$. Consequently, we repeat the same evaluation for Carso and compare the worst-case robustness amongst the two. In line with typical AT methods, and unlike diffusive purification, the robustness of Carso assessed by means of randAA is still lower w.r.t. than achieved by Pgd+EoT.
@@
Moving to scenario (b), Carso achieves a robust accuracy increase of $+27.47\%$ w.r.t. the classifier alone [Cui2023Decoupled], balanced by a $5.79\%$ decrease in clean accuracy. Our approach also improves upon the robust accuracy of the best AT-trained model [Wang2023Better] (WideResNet-70-16) by $23.98\%$. In the absence of a reliable robustness evaluation by means of Pgd+EoT for the best purification-based method [Lin2024Robust], we still obtain a $+20.25\%$ increase in robust accuracy upon its (largely overestimated) AA result.
@@
In scenario (c), Carso improves upon the classifier alone [Wang2023Better] (which is also the best AT-based approach for TinyImageNet-200) by $+22.26\%$. A significant clean accuracy toll is imposed by the relative complexity of the dataset, i.e. $-8.87\%$. In this setting, we lack any additional purification-based methods.
@@
Although the architecture of Carso is algorithmically differentiable end-to-end – and the integrated diagnostics of the randAA routines raised no warnings during the assessment – we additionally guard against the eventual gradient obfuscation [Athalye2018Obfuscated] induced by our method by repeating the evaluation at $\epsilon_{\infty}=0.95$, verifying that the resulting robust accuracy stays below random chance [Carlini2019Evaluating]. Results are shown in Table 3.
@@ 5.3 Limitations and open problems
In line with recent research aiming at the development of robust defences against multiple perturbations [Dolatabadi2021Robustness, Laidlaw2021Perceptual], our method determines a decrease in clean accuracy w.r.t. the original model on which it is built upon – especially in scenario (c) as the complexity of the dataset increases. This phenomenon is partly dependent on the choice of a VAE as the generative purification model, a requirement for the fairest evaluation possible in terms of robustness.
@@
Lastly, the scalability of Carso could be strongly improved by determining whether the internal representation used in conditional data generation may be restricted to a smaller subset of layers, while still maintaining the general robustness of the method.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
@@ Appendix A On Projected Gradient Descent adversarial training
The task of determining model parameters $\bm{\theta}^{\star}$ that are robust to adversarial perturbations is cast in [Madry2018Towards] as a min-max optimisation problem seeking to minimise adversarial risk, i.e.:
@@
The inner optimisation problem is solved, in [Madry2018Towards], by Projected Gradient Descent (Pgd), an iterative algorithm whose goal is the synthesis of an adversarial perturbation $\hat{\bm{\delta}}=\bm{\delta}^{\left(K\right)}$ after $K$ gradient ascent and projection steps defined as:
@@ Appendix B On the functioning of (conditional) Variational Autoencoders
Variational autoencoders (VAEs) [Kingma2014Auto, Rezende2014Stochastic] learn from data a generative distribution of the form $p(\bm{x},\bm{z})=p(\bm{x}\operatorname*{|}\bm{z})p(\bm{z})$, where the probability density $p(\bm{z})$ represents a prior over latent variable $\bm{z}$, and $p(\bm{x}\operatorname*{|}\bm{z})$ is the likelihood function, which can be used to sample data of interest $\bm{x}$, given $\bm{z}$.
@@
The setting is analogous in the case of conditional Variational Autoencoders [Sohn2015Learning, Yan2016Attribute2Image] (see section 3), where conditional sampling is achieved by
@@ Appendix C Justification of Adversarially-balanced batches
During the incipient phases of experimentation, preliminary tests were performed with the MNIST [LeCun2010MNIST] and Fashion-MNIST [Xiao2017Fashion] datasets – using a conditional VAE as the purifier, and small FCNs or convolutional ANNs as the classifiers. Adversarial examples were generated against the adversarially pre-trained classifier, and tentatively denoised by the purifier with one sample only. The resulting recovered inputs were classified by the classifier and the overall accuracy was recorded.
@@
Generating adversarial training examples by means of Pgd is considered the gold standard [Gowal2020Uncovering] and was first attempted as a natural choice to train the purifier. However, in this case, the following phenomena were observed:
@@
Appendix D Heuristic justification of the robust aggregation strategy

### v5

@@ 4.3 Hierarchical input and internal representation encoding
Training a conditional VAE requires (Sohn et al., 2015) that the conditioning set $\bm{c}$ is concatenated to the input $\bm{x}$ before encoding occurs, and to the sample of latent variables $\bm{z}$ right before decoding. The same is also true, with the suitable adjustments, for any conditional generative approach where the target and the conditioning set must be processed jointly.
In order to ensure the usability and scalability of Carso across the widest range of input data and classifier models, we propose to perform such processing in a hierarchical and partially disjoint fashion between the input and the conditioning set. In principle, the encoding of $\bm{x}$ and $\bm{c}$ can be performed by two different and independent subnetworks, until some form of joint processing must occur. This allows to retain the overall architectural structure of the purifier, while having finer-grained control over the inductive biases (Mitchell, 1980) deemed the most suitable for the respective variables.
@@
In order to use the VAE decoder at inference time, the entire compression machinery for the conditioning set must be preserved after training, and used to encode the internal representations extracted. The input encoder may be discarded instead.
@@ 4.5 Robust aggregation strategy
Such choice produces a robust prediction much harder to overtake in the event that an adversary selectively targets a specific input reconstruction. A heuristic analysis of this property (subsection D.1), together with an experimental justification for using such aggregation function within Carso (subsection D.2), are given in Appendix D.
@@ 5 Experimental assessment
Experimental evaluation of our method is carried out in terms of robust and clean image classification accuracy within three different scenarios (a, b, and c), determined by different classification tasks. The white-box threat model with a fixed $\ell_{\infty}$ norm bound is assumed throughout, as it generally constitutes the most demanding setup for adversarial defences.
@@ 5.1 Setup
The Cifar-10 (Krizhevsky, 2009) dataset is used in scenario (a), the Cifar-100 (Krizhevsky, 2009) dataset is used in scenario (b), whereas the TinyImageNet-200 (Chrabaszcz et al., 2017) dataset is used in scenario (c).
@@
In scenarios (a) and (b), the classifier is trained according to Cui et al. (2023); in scenario (c), according to Wang et al. (2023). Classifiers were always acquired as pre-trained models, using publicly available weights provided by the respective authors.
The purifier is trained on the VAE loss, using summed pixel-wise channel-wise binary cross-entropy as the reconstruction cost. Optimisation is performed by RAdam+Lookahead (Liu et al., 2020; Zhang et al., 2019b) with a learning rate schedule that presents a linear warm-up, a plateau phase, and a linear annealing (Smith, 2017). To promote the learning of meaningful reconstructions during the initial phases of training, the KL divergence term in the VAE loss is suppressed for an initial number of epochs. Afterwards, it is linearly modulated up to its actual value, during a fixed number of epochs ($\beta$ increase) (Higgins et al., 2017). The initial and final epochs of such modulation are reported in Table 15.
@@
In each scenario, we report the clean and robust test-set accuracy – the latter by means of AutoAttack (Croce & Hein, 2020a) – of the classifier and the corresponding Carso architecture.
For the classifier alone, the standard version of AutoAttack (AA) is used: i.e., the worst-case accuracy on a mixture of AutoPgd on the cross-entropy loss (Croce & Hein, 2020a) with $100$ steps, AutoPgd on the difference of logits ratio loss (Croce & Hein, 2020a) with $100$ steps, Fab (Croce & Hein, 2020b) with $100$ steps, and the black-box Square attack (Andriushchenko et al., 2020) with $5000$ queries.
In the evaluation of the Carso architecture, the number of reconstructed samples per input is set to $8$, the logits are aggregated as explained in subsection 4.5, and the output class is finally selected as the $\operatorname*{\arg\max}$ of the aggregation. Due to the stochastic nature of the purifier, robust accuracy is assessed by a version of AutoAttack suitable for stochastic defences (randAA) – composed of AutoPgd on the cross-entropy and difference of logits ratio losses, across $20$ Expectation over Transformation (EoT) (Athalye et al., 2018b) iterations with $100$ gradient ascent steps each. In the specific case of scenario (a), we also assess our method by the Pgd+EoT pipeline proposed by Lin et al. (2024), as explained in subsection 5.2.
@@
All experiments were performed on an NVIDIA DGX A100 system. Training in scenarios (a) and (c) was run on 8 NVIDIA A100 GPUs with 40 GB of dedicated memory each; in scenario (b) 4 of such devices were used. Elapsed training time for the purifier in all scenarios is reported in Table 1.
Table 1: Elapsed running time for training the purifier in the different scenarios considered.
@@ 5.2 Results and discussion
Table 2: Clean (results in italic) and adversarial (results in upright) accuracy for the different models and datasets used in the respective scenarios. The following abbreviations are used: Scen: scenario considered; AT/Cl: clean accuracy for the adversarially-pretrained model used as the classifier, when considered alone; C/Cl: clean accuracy for the Carso architecture; AT/AA: robust accuracy (by the means of AutoAttack) for the adversarially-pretrained model used as the classifier, when considered alone; C/randAA: robust accuracy for the Carso architecture, when attacked end-to-end by AutoAttack for randomised defences; Best AT/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial training alone (any model); Best P/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial purification (any model). Robust accuracies in round brackets are obtained using the Pgd+EoT (Lee & Kim, 2024) pipeline, developed for diffusion-based purifiers. The best clean and robust accuracies per dataset are shown in bold. The clean accuracies for the models referred to in the Best columns are shown in Table 16 (in Appendix F).
Scen. Dataset AT/Cl C/Cl AT/AA C/randAA (Pgd+EoT) Best AT/AA Best P/AA (Pgd+EoT) (a) Cifar-10 0.9216 0.8686 0.6773 0.7613 (0.7689) 0.7371 0.7812 (0.6641) (b) Cifar-100 0.7385 0.6806 0.3918 0.6665 0.4267 0.4609 (c) TinyImageNet-200 0.6519 0.5632 0.3130 0.5356 0.3130
@@
Comparing the robust accuracy of the classifier model used in scenario (a) (Cui et al., 2023) with that resulting from the inclusion of the same model in the Carso architecture, we observe a $+8.4\%$ increase. This is counterbalanced by a $-5.6\%$ clean accuracy toll. The same version of Carso further provides a $+2.42$ robustness increase w.r.t. the current best AT-trained model (Bartoldson et al., 2024) that employs a 4$\sim$ 4$\times$ 4 × larger WideResNet-96-16 model.
In addition, our method provides a remarkable $+9.72\%$ increase in robust accuracy w.r.t. to the best adversarial purification approach (Lin et al., 2024), a diffusion-based purifier. However, the comparison is not as straightforward. In fact, the original paper (Lin et al., 2024) reports a robust accuracy of $78.12\%$ using AutoAttack on the gradients obtained via the adjoint method (Nie et al., 2022). As noted in Lee & Kim (2024), such evaluation (which uses the version of AutoAttack that is unsuitable for stochastic defences) leads to a large overestimation of the robustness of diffusive purifiers. As suggested in Lee & Kim (2024), the authors of Lin et al. (2024) re-evaluate the robust accuracy according to a more suitable pipeline (Pgd+EoT, whose hyperparameters are shown in Table 13), obtaining a much lower robust accuracy of $66.41\%$. Consequently, we repeat the same evaluation for Carso and compare the worst-case robustness amongst the two. In line with typical AT methods, and unlike diffusive purification, the robustness of Carso assessed by means of randAA remains lower w.r.t. that achieved by Pgd+EoT.
@@
Moving to scenario (b), Carso achieves a robust accuracy increase of $+27.47\%$ w.r.t. the classifier alone (Cui et al., 2023), balanced by a $-5.79\%$ decrease in clean accuracy. Our approach also improves upon the robust accuracy of the best AT-trained model (Wang et al., 2023) (WideResNet-70-16) by $+23.98\%$. In the absence of a reliable robustness evaluation by means of Pgd+EoT for the best purification-based method (Lin et al., 2024), we still obtain a $+20.25\%$ increase in robust accuracy upon its (largely overestimated) AA result.
@@
In scenario (c), Carso improves upon the classifier alone (Wang et al., 2023) (which is also the best AT-based approach for TinyImageNet-200) by $+22.26\%$. A significant clean accuracy toll is imposed by the relative complexity of the dataset, i.e. $-8.87\%$. In this setting, we lack any additional purification-based methods.
@@
Although the architecture of Carso is algorithmically differentiable end-to-end – and the integrated diagnostics of the randAA routines raised no warnings during the assessment – we additionally guard against the eventual gradient obfuscation (Athalye et al., 2018a) induced by our method by repeating the evaluation at $\epsilon_{\infty}=0.95$, verifying that the resulting robust accuracy stays below random chance (Carlini et al., 2019). Results are shown in Table 3.
@@ 5.3 Limitations and open problems
In line with recent research aiming at the development of robust defences against multiple perturbations (Dolatabadi et al., 2022; Laidlaw et al., 2021), our method produces a decrease in clean accuracy w.r.t. the original model on which it is built upon – especially in scenario (c) as the complexity of the classification task increases. This phenomenon is partially dependent on the choice of a VAE as the generative purification model, a requirement for the fairest evaluation possible in terms of robustness.
@@
Additionally, the scalability of Carso could be strongly improved by determining whether the internal representation used in conditional data generation may be restricted to a smaller subset of layers, while still maintaining the general robustness of the method.
Finally, a thorough investigation of the normalised doubly-exponential logit product aggregation strategy needs to be undertaken in order to shed some light on the specific mechanisms that lead to the much improved defensive capabilities of the system.
@@ 6 Conclusion
Broader Impact Statement
In this work, we investigate the use of a novel technique for the improvement of adversarial robustness in image classification models. Our method does not carry the potential for additional risks or implications in sensitive areas, in comparison to the original models used. Indeed, the goal of our technique is instead to ultimately improve and enhance the robustness of classification models to make them more trustworthy, predictable, and resistant to tampering.
Acknowledgements
@@ Appendix A On Projected Gradient Descent adversarial training
The task of determining model parameters $\bm{\theta}^{\star}$ that are robust to adversarial perturbations is cast in (Madry et al., 2018) as a min-max optimisation problem seeking to minimise adversarial risk, i.e.:
@@
The inner optimisation problem is solved, by Madry et al. (2018), by Projected Gradient Descent (Pgd), an iterative algorithm whose goal is the synthesis of an adversarial perturbation $\hat{\bm{\delta}}=\bm{\delta}^{\left(K\right)}$ after $K$ gradient ascent and projection steps defined as:
@@ Appendix B On the functioning of (conditional) Variational Autoencoders
Variational autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) learn from data a generative distribution of the form $p(\bm{x},\bm{z})=p(\bm{x}\operatorname*{|}\bm{z})p(\bm{z})$, where the probability density $p(\bm{z})$ represents a prior over latent variable $\bm{z}$, and $p(\bm{x}\operatorname*{|}\bm{z})$ is the likelihood function, which can be used to sample data of interest $\bm{x}$, given $\bm{z}$.
@@
The setting is analogous in the case of conditional Variational Autoencoders (Sohn et al., 2015; Yan et al., 2016) (see section 3), where conditional sampling is achieved by
@@ Appendix C Justification of Adversarially-balanced batches
During the incipient phases of experimentation, preliminary tests were performed with the MNIST (LeCun & Cortes, 2010) and Fashion-MNIST (Xiao et al., 2017) datasets – using a conditional VAE as the purifier, and small FCNs or convolutional ANNs as the classifiers. Adversarial examples were generated against the adversarially pre-trained classifier, and tentatively denoised by the purifier with one sample only. The resulting recovered inputs were classified by the classifier and the overall accuracy was recorded.
@@
Generating adversarial training examples by means of Pgd is considered the gold standard (Gowal et al., 2020) and was first attempted as a natural choice to train the purifier. However, in this case, the following phenomena were observed:
@@
Appendix D Justification of the robust aggregation strategy
@@ D.1 Heuristic analysis
D.2 Experimental analysis
To further corroborate the choice of the specific aggregation function described in subsection 4.5, we repeat the assessment of Carso under the same conditions described in section 5, the only difference being the use of the alternative aggregation functions analysed in subsection D.1 (i.e. logit and probability aggregation). Results, in terms of both clean and robust accuracy, are shown in Table 4.
As we can see, the use of the alternative aggregation strategies leads to a minimal variations in the clean accuracy attained, whereas the corresponding robust accuracy sharply decreases – in the case of Cifar-10 even below random chance – as the attacks become increasingly effective. Such results strongly corroborate the use of the normalised doubly-exponential logit product proposed as the aggregation strategy in subsection 4.5 and prove its central role in the overall adversarial robustness of the method.
Table 4: Clean (results in italic) and adversarial (results in upright) accuracy for alternative aggregation strategies used within Carso. The following abbreviations are used: Scen: scenario considered; C/Cl (L.A.): clean accuracy of Carso with logit average aggregation; C/randAA (L.A.): robust accuracy of Carso with logit average aggregation, assessed by means of AutoAttack for stochastic defences; C/Cl (P.A.): clean accuracy of Carso with probability average aggregation; C/randAA (P.A.): robust accuracy of Carso with probability average aggregation, assessed by means of AutoAttack for stochastic defences; C/Cl: clean accuracy of Carso with our proposed aggregation; C/randAA: robust accuracy of Carso with our proposed aggregation, assessed by means of AutoAttack for stochastic defences. Results from the last two columns mirror those of Table 2.
Scen. Dataset C/Cl (L.A.) C/randAA (L.A.) C/Cl (P.A.) C/randAA (P.A.) C/Cl C/randAA (a) Cifar-10 0.8688 0.0086 0.8688 0.0092 0.8686 0.7613 (b) Cifar-100 0.6808 0.0436 0.6807 0.0439 0.6806 0.6665